# A Simple but Strong Baseline for Sounding Video Generation: Effective Adaptation of Audio and Video

### v1

@@ Abstract
1 Introduction
Diffusion models have made great strides in the last few years in various generation tasks across modalities including image, video, and audio (Yang et al., 2023). Although these models are often large-scale and require a huge amount of computational resources for training, several prior studies such as Stable Diffusion (Rombach et al., 2022) have made their trained models publicly available, which substantially accelerates the progress of research and development on generative models. However, these models have mainly focused on a single modality, and it is still challenging to construct a model that is capable of generating multi-modal data.
In this work, we focus on audio-video joint generation, which is also known as sounding video generation (Liu et al., 2023b). Although sounding videos are one of the most popular types of multi-modal data, their generation has been addressed by only a few recent studies (Liu et al., 2023b; Ruan et al., 2023; Tang et al., 2023) due to the extremely high difficulty of handling heterogeneous and high-dimensional data for generative modelling. This challenge makes the training of multi-modal generative models much more expensive than that of single-modal models, and it creates a barrier to the research and development of sounding-video generation technologies.
@@ 1 Introduction
2 Background and related work
2.1 Diffusion models
Diffusion models (Yang et al., 2023) are a family of generative models designed to generate data by reversing a diffusion process. Here, we briefly review one of the most popular types of diffusion models, called the denoising diffusion probabilistic model (Ho et al., 2020).
2.1.1 Basics
@@
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),\ i.e.\ \mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (2)
@@
$\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\ \sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},$ (4)
where $\epsilon_{\theta}$ represents the model with learnable parameters $\theta$ for the noise prediction. It is also well-known that we can sample $\mathbf{x}_{t-1}$ in a deterministic manner using DDIM (Song et al., 2020), as:
@@
2.1.2 Application to single-modal generation
@@
Diffusion models have demonstrated remarkable performance across various modalities, particularly in vision and audio domains. In the vision domain, the initial attempt was limited to generating low-resolution images (Ho et al., 2020), but it was soon extended to handle high-resolution images (Rombach et al., 2022; Saharia et al., 2022) and videos (Ho et al., 2022; Guo et al., 2023; Blattmann et al., 2023). To reduce the computational cost due to the high dimensionality of data, the latest diffusion models are often trained in the space of latent features (Rombach et al., 2022) obtained by an encoder such as VAE (Kingma & Welling, 2014) or VQGAN (Esser et al., 2021). A similar trend can be found in the audio domain: diffusion models were initially used to directly generate waveforms (Chen et al., 2020b; Kong et al., 2020) and were then extended to generate compressed representation or latent features of audio signals (Liu et al., 2023a; Huang et al., 2023). In this work, we utilize the latest publicly available models in both domains, specifically, AnimateDiff (Guo et al., 2023) and AudioLDM (Liu et al., 2023a), to efficiently construct an audio-visual generative model that is capable of jointly generating video and audio well aligned with each other.
2.2 Audio-video generative models
2.2.1 Cross-modal conditional generation
@@
Video-conditioned audio generation (V2A) has been extensively explored in the literature of audio-visual generative models. Pioneering works adopted regression models (Owens et al., 2016; Chen et al., 2020c; Ghose & Prevost, 2020) and GANs (Hao et al., 2018; Ghose & Prevost, 2022), but auto-regressive models (Iashin & Rahtu, 2021; Du et al., 2023) and diffusion models (Luo et al., 2023; Mo et al., 2023; Su et al., 2023; Comunità et al., 2024; Wang et al., 2024) have become popular choices recently due to their scalability and capability of generating diverse data. To apply these models for V2A, we additionally need a mechanism to feed video conditional information into audio generation models. Such cross-modal conditioning has typically been achieved by a cross-attention mechanism (Vaswani et al., 2017), where the conditional information is used to compute keys and values in the attention process. In this work, we propose a new module for the cross-modal conditioning that is simple but effective for obtaining higher alignment between modalities.
@@
Compared with V2A, audio-conditioned video generation (A2V) has not been as intensely addressed in the literature, as high-quality video generation itself is already challenging. Given the success of large-scale autoregressive models (Weissenborn et al., 2020; Yan et al., 2021) and diffusion models (Ho et al., 2022; Guo et al., 2023; Blattmann et al., 2023) in video generation, audio-to-video generation has also been addressed by extending these models to accept audio conditions (Ge et al., 2022; Yariv et al., 2023; Zhang et al., 2024a). In this paper, we also extend existing diffusion models of video generation, but our goal is to enable joint generation of audio and video, which is substantially more challenging than audio-to-video. For this purpose, we propose a new mechanism to adjust timesteps across modalities during the generation process. It is particularly required for joint generation to effectively handle noisy multi-modal data at each timestep for higher alignment, because any clean data is not accessible during the generation process differently from the situation of V2A or A2V.
2.2.2 Audio-video joint generation
@@
As mentioned above, audio-video joint generation, namely, sounding video generation, is challenging compared to single-modal generation, and few studies have tackled it (Liu et al., 2023b; Ruan et al., 2023; Tang et al., 2023). SVG-VQGAN (Liu et al., 2023b) adopts a novel tokenizer for audio and video to obtain suitable representation for multi-modal generation with auto-regressive models. MM-Diffusion (Ruan et al., 2023) and TAVDiffusion (Mao et al., 2024) are multi-modal diffusion models specifically designed for audio-video paired data. CoDi (Tang et al., 2023) integrates several single-modal dedicated diffusion models and additionally adopts environment encoders to extract modality-specific features to condition the generation process in the other modalities. All these models incur a large computational cost for training due to their new model architectures specialized for the joint generation. In this work, we aim to construct sounding video generation models with minimal effort by effectively transferring state-of-the-art models in both the audio and video domains. A very recent work by Xing et al. (2024) shares a similar motivation to ours, but it adopts guidance based approach, which heavily limits the capability of the model to generate temporally aligned samples. In contrast, we introduce an efficient adaptation method that significantly enhances temporal alignment across modalities.
3 Proposed method
@@ 3.1 Overview
Our goal is to build a single model capable of generating video and audio jointly, utilizing two pre-trained diffusion models, one for video and the other for audio. These models, referred to as base models, are each represented by a U-Net structured neural network with pre-trained parameters. Our model, as shown in Figure 1, includes two U-Nets as base models with pre-trained modules depicted by gray rectangles. To enable joint generation of aligned video and audio, self-attention blocks are inserted into each U-Net, and connectors are introduced to extract features at each modality. These features are then fed into the U-Net of the other modality, allowing the model to utilize all modal information for noise prediction, resulting in better alignment across modalities, which will be described in Section 3.3. Note that, in the experiments, the noise prediction is conditioned by a given text prompt, as we used text-conditional generative models as the base ones. The text condition is fed into each U-Net in the same way as the original base models, and the same text prompt is used for both modalities.
During training, only the newly introduced modules are updated, while the pre-trained modules of each U-Net remain fixed. Like standard latent diffusion models, our model predicts noise from a pair of noisy latents and outputs slightly denoised latents. A key difference lies in the timestep setting, where different timesteps are set for each modality. This is due to the original design of the timestep in the U-Net at each modality, which may not be suitable for multi-modal joint generation. This issue and our solution are discussed in more detail in Section 3.2.
3.2 Timestep adjustment
3.2.1 Why do we need to adjust timesteps across modalities?
@@
To visualize this discrepancy, we plot the loss distribution over the timestep in Fig. 2. The loss values are measured in the experiment (described in Section 4.1) and are normalized by their value at $t=0$ for each modality. Here, we choose the loss instead of signal-to-noise ratio (SNR) as a proxy to observe how samples are generated, as SNR is not suitable for comparisons between data with a different number of dimensions (Hoogeboom et al., 2023). Obviously, the loss on video data is heavily skewed towards $t=0$, which implies that the noise schedule for videos is set to more rapidly collapse data into noise as the timestep progresses. Such a noise schedule is often adopted to address the high dimensionality of video data (Hoogeboom et al., 2023). However, if we directly re-use this schedule in the joint generation setting, video information will not be very informative for audio generation at the intermediate timesteps, which makes the generation process more like audio-to-video than joint generation. To solve this problem, we need to adjust the timesteps across modalities.
(a) Without timestep adjustment
(b) With timestep adjustment
@@
3.2.2 A simple solution for timestep adjustment
@@
Figure 2 shows the loss distributions after applying the adjustment with $\gamma=1.5$. The horizontal axis represents the global timestep, and the vertical axis represents the normalized loss at each local timestep corresponding to the global one. Compared with Fig. 2, the loss distributions become much more similar to each other. This indicates that how samples are generated along with the timestep becomes more aligned between video and audio. Consequently, through the joint generation, we can expect higher alignment between the generated pair of data. In this paper, we set $\gamma$ to $1.5$, unless otherwise noted. How to automatically set this hyperparameter remains as future work.
3.3 How to feed cross-modal features into U-Net
3.3.1 The standard choice: cross-attention and its limitation
In the literature, the cross-attention mechanism has been extensively used for cross-modal conditioning in diffusion models. Figure 3 shows the simplest design in this approach adopted in our baseline method (Tang et al., 2023). For brevity, we discuss the case of audio-conditioned video generation, but it can also be applied to the case of video-conditioned audio generation. In this design, the conditional audio information is embedded into a single feature vector by an encoder, and it is used to compute keys and values in the cross-attention taken with the intermediate features in the video generation model. By training the encoder and the cross-attention block with audio-video paired data, we can make the model generate videos aligned with given audio information. Although this design is simple and widely applicable, it is quite challenging to achieve higher temporal consistency between the audio condition and generated video, since the single vector does not have sufficient capability to represent every piece of temporally local information in the audio condition.
To boost the capability of the embedding features, we can extend the above-mentioned design by using multiple vectors each of which represents the temporally local information of conditional audio as done in Yariv et al. (2023). However, we still cannot strongly guarantee the temporal alignment, as it provides too much flexibility to connect the temporally-local audio information with the video to be generated, which may cause mis-alignment. It is also possible to adopt a more sophisticated attention mechanism (Ruan et al., 2023) or specifically dedicated encoder for embedding (Luo et al., 2023), but this substantially reduces applicability to the existing audio and video generation models, which does not fit our goal in this work.
@@ 3.3.1 The standard choice: cross-attention and its limitation
3.3.2 Cross-Modal Conditioning as Positional Encoding (CMC-PE)
To achieve higher temporal alignment, we propose Cross-Modal Conditioning as Positional Encoding (CMC-PE), a simple but effective method of cross-modal conditioning. Figure 3 depicts how CMC-PE works. First, the conditional audio is encoded to a sequence of feature vectors along with time frames that work as if representing temporal position information. The extracted features are then added to the intermediate features in the video U-Net to function as positional embedding. To make this addition process valid, the features are interpolated and broadcast in advance to match their shape with that of the video features. Finally, the updated features are processed with a self-attention block. The features used for CMC-PE are extracted from current noisy latents $\mathbf{x}_{t}$ by the connector. We adopt the self-conditioning technique here (Chen et al., 2022), where the estimated data $\hat{\mathbf{x}}_{0|t}$ at each timestep shown in Eq. (6) is concatenated to the input.
@@ 3.4 Training and inference
where $s\in[\mathrm{v},\mathrm{a}]$ indicates the modality where the loss is to be computed, and $t_{\mathrm{v}}$ and $t_{\mathrm{a}}$ represent the local timestep for video and audio, respectively. A key point here is that the local timesteps are independently sampled from a uniform distribution. Due to this, the loss is computed on all possible combinations of the local timesteps so that the trained model can handle the timestep adjustment with any value of $\gamma$ specified in the inference phase. During the training of the model, the connectors and the inserted self-attention blocks are optimized with audio-video paired data, while the pre-trained parameters of the other modules are fixed, with one small exception (discussed in the appendix A.2).
@@
The generation process in our method is almost the same as that in usual diffusion models except for the setting of local timesteps. In each step of the generation process, the model predicts noise for video and audio latents following the local timestep setting. Once these noises are predicted, we can apply any inference technique, also called a “sampler” (Yang et al., 2023), to obtain $\mathbf{x}_{m(t-1)}^{(\mathrm{v})}$ and $\mathbf{x}_{n(t-1)}^{(\mathrm{a})}$ in the same manner as in the base models. In this paper, we use one of the most popular ones, DDIM (Song et al., 2020) shown in Eq. (5), for both modalities. The generation process in the proposed method is summarized in the appendix A.1.
4 Experiments
4.1 Experiments with a dedicated dataset for evaluating temporal alignment
4.1.1 Dataset and evaluation metrics
We extended the GreatestHits dataset (Owens et al., 2016) for our experiments. It contains 977 videos of humans hitting various objects with a drumstick in the scene. As the hitting sound and motion are dominant in the video, this dataset is suitable for evaluating the temporal alignment between the generated video and audio. We created video captions using LLaVA-NeXT (Zhang et al., 2024b) and utilized them as text conditions in our method. The details of this process are described in the appendix A.3.
We evaluated the quality of the generated data from three perspectives: video quality, audio quality, and temporal alignment. For the former two, we used FVD (Unterthiner et al., 2018) and FAD (Kilgour et al., 2019), both of which are widely utilized in the literature. To measure how much the generated video and audio are aligned with each other in term of temporal dynamics, we used the AV-Align score proposed by Yariv et al. (2023). This score is defined as Intersection-over-Union between onsets detected from the audio and peaks obtained from the optical flow. It is especially useful to measure the temporal alignment in the GreatestHits dataset, as hitting sounds make clear onsets, and hitting motions have correlating and distinct peaks in the optical flow.
We slightly modified how to compute AV-Align score from the official implementation. Specifically, we tuned hyper-parameters of the optical flow estimation and those of the onset detection to accurately estimate hitting timing using annotated timestamps in the Greatest Hits dataset. In addition, we compute IoU after rewriting it with precision and recall to mitigate an issue caused by the difference of temporal resolution between video and audio. Details are provided in the appendix A.4.
4.1.2 Setup
  1. 1.
    One with the same setting as CoDi (Tang et al., 2023), in which cross-attention blocks are used for cross-modal conditioning.
  2. 2.
  3. 3.
For the training, we used the Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1e-5, and the batch size and the number of epochs were set to 16 and 1,000, respectively. For generation, we set the number of global timesteps $T$ to 25. We adopted classifier-free guidance (Ho & Salimans, 2021) at each modality and set the strength of the guidance to 7.5 and 2.5 for video and audio, respectively, which are the standard settings in the original base models.
4.1.3 Results
Table 1 shows the evaluation results. Replacing cross-attention with CMC-PE improves the AV-Align score as well as FVD and FAD, which demonstrates that CMC-PE has a better inductive bias for temporal alignment than the cross-attention mechanism. The AV-Align score is further improved by conducting the timestep adjustment when generating data. This is achieved by making the generation process in both modalities mutually informative as discussed in Section 3.2. Meanwhile, using too large a $\gamma$ leads to degradation of the performance due to the deviation from the original noise schedule. Overall, the proposed method performs substantially better in terms of the cross-modal alignment than our baseline following the design in CoDi (Tang et al., 2023).
Table 1: Experimental results with the GreatestHits dataset.
Method FVD ($\downarrow$) FAD ($\downarrow$) AV-Align ($\uparrow$)
Cross-attention (same as CoDi (Tang et al., 2023)) 379 2.35 0.250
Our method without timestep adjustment ($\gamma=1$) 393 1.29 0.256
Our method ($\gamma=1.25$) 387 1.32 0.257
Our method ($\gamma=1.50$) 381 0.60 0.268
Our method ($\gamma=1.75$) 374 0.61 0.268
Our method ($\gamma=2.00$) 383 1.32 0.265
4.2 Experiments with benchmark datasets
4.2.1 Dataset and evaluation metrics
To compare the proposed method with existing methods, we conducted experiments with two popular benchmark datasets: Landscape (Lee et al., 2022) and VGGSound (Chen et al., 2020a). The Landscape dataset consists of 928 videos covering nine classes of natural scenes, while VGGSound is a substantially larger and more diverse dataset containing nearly 200K video clips covering about 300 sound classes. To enhance the data quality, we filtered 60K videos in which audio-video alignment is weak, as done in TempoToken (Yariv et al., 2023). In both datasets, we used the class names as the text conditions and trained our model to generate four-second audio-video pairs. The video comprises four frames per second, and the size of each frame is 256 $\times$ 256. The sampling rate of the audio is 16 kHz. Note that we changed fps from the previous experiments to follow the setting in the prior studies (Iashin & Rahtu, 2021; Luo et al., 2023; Yariv et al., 2023).
Differently from the previous experiments, we did not use AV-Align for the evaluation, as the videos in both Landscape and VGGSound often lack distinct motions highly correlating their audios. Instead, we used ImageBind score (Girdhar et al., 2023) between audio and video (IB-AV) to evaluate the cross-modal semantic alignment. Additionally, we computed the ImageBind score for text-audio and text-video pairs (IB-TA and IB-TV, respectively) to evaluate the audio and video quality in terms of fidelity to text condition.
4.2.2 Setup
T2A2V
We used TempoToken (Yariv et al., 2023) to re-generate videos from the audios that are generated by the proposed method.
T2V2A
We used SpecVQGAN (Iashin & Rahtu, 2021) and DiffFoley (Luo et al., 2023) to re-generate audios from the videos that are generated by the proposed method.
Joint generation
We used MM-Diffusion (Ruan et al., 2023). For a fair comparison, the number of timesteps was set to be the same as that of the proposed method.
For all methods, we used the official implementation and pretrained models provided by the respective authors. Note that the pretrained models of SpecVQGAN and DiffFoley were available for VGGSound, and that of MM-Diffusion was available for Landscape. When we cannot specify frame rate or resolution of the generated data, we resized the generated data to make it match our setting before the evaluation. For the training of our model, the batch size and the number of epochs were set to 16 and 100 for the Landscape dataset, and to 128 and 30 for VGGSound, respectively. The other settings are the same as those in the previous experiments.
4.2.3 Results
Tables 2 and 3 show the results with Landscape and VGGSound, respectively. Firstly, TempoToken failed to generate high-quality videos, which indicates that it is not robust against even small artifacts in the conditional audio caused by preceding text-to-audio generation. This is one of the most critical issues of sequential approaches like T2A2V and T2V2A, and SpecVQGAN and DiffFoley also suffered from it, resulting in relatively low audio quality. In contrast, the proposed method achieved the best quality in both video and audio except for FVD in Landscape and IB-AV in VGGSound while attaining high cross-modal alignment. This indicates the importance of the training dedicated to joint generation and the effectiveness of our method.
Table 2: Experimental results with Landscape dataset.
Method FVD ($\downarrow$) IB-TV ($\uparrow$) FAD ($\downarrow$) IB-TA ($\uparrow$) IB-AV ($\uparrow$)
TempoToken (Yariv et al., 2023) $>$ 3000 0.220 0.146
MM-Diffusion (Ruan et al., 2023) 1689 16.4 0.191
Proposed method 1122 0.238 6.63 0.146 0.192
Table 3: Experimental results with VGGSound dataset. ($\dagger$ DiffFoley uses a larger dataset for learning cross-modal alignment)
Method FVD ($\downarrow$) IB-TV ($\uparrow$) FAD ($\downarrow$) IB-TA ($\uparrow$) IB-AV ($\uparrow$)
TempoToken (Yariv et al., 2023) 2473 0.155 0.168
SpecVQGAN (Iashin & Rahtu, 2021) 5.08 0.059 0.100
DiffFoley (Luo et al., 2023) 5.72 0.074 0.159
Proposed method 333 0.277 1.46 0.129 0.155
4.3 Limitation
In the experiments, we observed that our model occasionally ignores some parts of the textual condition. For example, in Fig. 4, while the conditional text contains “a blue plastic trash bag,” the word blue is somewhat ignored in the generated video. We conjecture that this is caused by focusing too much on aligning with the audio to be jointly generated that do not contain visual information (e.g. color). How to simultaneously achieve fine-grained cross-modal alignment and semantic alignment with the textual condition would be an interesting avenue for future work. Additionally, as our method leverages pre-trained models, its performance heavily depends on that of these base models.
5 Conclusion
Appendix A Appendix
@@
A.1 Generation process in the proposeed method
The generation process in the proposed method is summarized in Algorithm 1.
A.2 Implementation details

### v3

@@ Abstract
Index Terms:
multi-modal generation, diffusion models, audio-visual.
I Introduction
Diffusion models have made great strides in the last few years in various generation tasks across modalities including image, video, and audio [1]. Although these models are often large-scale and require a huge amount of computational resources for training, several prior studies such as Stable Diffusion [2] have made their trained models publicly available, which substantially accelerates the progress of research and development on generative models. However, these models have mainly focused on a single modality, and it is still challenging to construct a model that is capable of generating multi-modal data.
In this work, we focus on audio-video joint generation, which is also known as sounding video generation [3]. Although sounding videos are one of the most popular types of multi-modal data, their generation has been addressed by only a few recent studies [3, 4, 5] due to the extremely high difficulty of handling heterogeneous and high-dimensional data for generative modelling. This challenge makes the training of multi-modal generative models much more expensive than that of single-modal models, and it creates a barrier to the research and development of sounding-video generation technologies.
@@ I Introduction
II Background and related work
II-A Diffusion models
Diffusion models [1] are a family of generative models designed to generate data by reversing a diffusion process. Here, we briefly review one of the most popular types of diffusion models, called the denoising diffusion probabilistic model [6].
II-A1 Basics
@@
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (2)
$\ i.e.\ \mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$
@@
$\mu_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),$ (4)
$\sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},$ (5)
where $\epsilon_{\theta}$ represents the model with learnable parameters $\theta$ for the noise prediction. It is also well-known that we can sample $\mathbf{x}_{t-1}$ in a deterministic manner using DDIM [7], as:
@@
II-A2 Application to single-modal generation
@@
Diffusion models have demonstrated remarkable performance across various modalities, particularly in vision and audio domains. In the vision domain, the initial attempt was limited to generating low-resolution images [6], but it was soon extended to handle high-resolution images [2, 8] and videos [9, 10, 11]. To reduce the computational cost due to the high dimensionality of data, the latest diffusion models are often trained in the space of latent features [2] obtained by an encoder such as VAE [12] or VQGAN [13]. A similar trend can be found in the audio domain: diffusion models were initially used to directly generate waveforms [14, 15] and were then extended to generate compressed representation or latent features of audio signals [16, 17]. In this work, we utilize the latest publicly available models in both domains, specifically, AnimateDiff [10] and AudioLDM [16], to efficiently construct an audio-visual generative model that is capable of jointly generating video and audio well aligned with each other.
II-B Audio-video generative models
II-B1 Cross-modal conditional generation
@@
Video-conditioned audio generation (V2A) has been extensively explored in the literature of audio-visual generative models. Pioneering works adopted regression models [18, 19, 20] and GANs [21, 22], but auto-regressive models [23, 24] and diffusion models [25, 26, 27, 28, 29, 30] have become popular choices recently due to their scalability and capability of generating diverse data. To apply these models for V2A, we additionally need a mechanism to feed video conditional information into audio generation models. Such cross-modal conditioning has typically been achieved by a cross-attention mechanism [31], where the conditional information is used to compute keys and values in the attention process. In this work, we propose a new module for the cross-modal conditioning that is simple but effective for obtaining higher alignment between modalities.
@@
Compared with V2A, audio-conditioned video generation (A2V) has not been as intensely addressed in the literature, as high-quality video generation itself is already challenging. Given the success of large-scale autoregressive models [32, 33] and diffusion models [9, 10, 11] in video generation, audio-to-video generation has also been addressed by extending these models to accept audio conditions [34, 35, 36]. In this paper, we also extend existing diffusion models of video generation, but our goal is to enable joint generation of audio and video, which is substantially more challenging than audio-to-video. For this purpose, we propose a new mechanism to adjust timesteps across modalities during the generation process. It is particularly required for joint generation to effectively handle noisy multi-modal data at each timestep for higher alignment, because any clean data is not accessible during the generation process differently from the situation of V2A or A2V.
II-B2 Audio-video joint generation
@@
As mentioned above, audio-video joint generation, namely, sounding video generation, is challenging compared to single-modal generation, and few studies have tackled it [3, 4, 5]. SVG-VQGAN [3] adopts a novel tokenizer for audio and video to obtain suitable representation for multi-modal generation with auto-regressive models. MM-Diffusion [4], TAVDiffusion [37], and AVDiT [38] are multi-modal diffusion models specifically designed for audio-video paired data. CoDi [5] integrates several single-modal dedicated diffusion models and additionally adopts environment encoders to extract modality-specific features to condition the generation process in the other modalities. All these models incur a large computational cost for training due to their new model architectures specialized for the joint generation. In this work, we aim to construct sounding video generation models with minimal effort by effectively transferring state-of-the-art models in both the audio and video domains. A very recent work by [39] shares a similar motivation to ours, but it adopts guidance based approach, which heavily limits the capability of the model to generate temporally aligned samples. In contrast, we introduce an efficient adaptation method that significantly enhances temporal alignment across modalities.
III Proposed method
@@ III-A Overview
Our goal is to build a single model capable of generating video and audio jointly, utilizing two pre-trained diffusion models, one for video and the other for audio. These models, referred to as base models, are each represented by a U-Net structured neural network with pre-trained parameters. Our model, as shown in Figure 1, includes two U-Nets as base models with pre-trained modules depicted by gray rectangles. To enable joint generation of aligned video and audio, self-attention blocks are inserted into each U-Net, and connectors are introduced to extract features at each modality. These features are then fed into the U-Net of the other modality, allowing the model to utilize all modal information for noise prediction, resulting in better alignment across modalities, which will be described in Section III-C. Note that, in the experiments, the noise prediction is conditioned by a given text prompt, as we used text-conditional generative models as the base ones. The text condition is fed into each U-Net in the same way as the original base models, and the same text prompt is used for both modalities.
During training, only the newly introduced modules are updated, while the pre-trained modules of each U-Net remain fixed. Like standard latent diffusion models, our model predicts noise from a pair of noisy latents and outputs slightly denoised latents. A key difference lies in the timestep setting, where different timesteps are set for each modality. This is due to the original design of the timestep in the U-Net at each modality, which may not be suitable for multi-modal joint generation. This issue and our solution are discussed in more detail in Section III-B.
III-B Timestep adjustment
III-B1 Why do we need to adjust timesteps across modalities?
@@
To visualize this discrepancy, we plot the loss distribution over the timestep in Fig. 2. The loss values are measured in the experiment (described in Section IV-A) and are normalized by their value at $t=0$ for each modality. Here, we choose the loss instead of signal-to-noise ratio (SNR) as a proxy to observe how samples are generated, as SNR is not suitable for comparisons between data with a different number of dimensions [40]. Obviously, the loss on video data is heavily skewed towards $t=0$, which implies that the noise schedule for videos is set to more rapidly collapse data into noise as the timestep progresses. Such a noise schedule is often adopted to address the high dimensionality of video data [40]. However, if we directly re-use this schedule in the joint generation setting, video information will not be very informative for audio generation at the intermediate timesteps, which makes the generation process more like audio-to-video than joint generation. To solve this problem, we need to adjust the timesteps across modalities.
(a) w/o timestep adjustment
(b) w/ timestep adjustment
@@
III-B2 A simple solution for timestep adjustment
@@
Figure 2 shows the loss distributions after applying the adjustment with $\gamma=1.5$. The horizontal axis represents the global timestep, and the vertical axis represents the normalized loss at each local timestep corresponding to the global one. Compared with Fig. 2, the loss distributions become much more similar to each other. This indicates that how samples are generated along with the timestep becomes more aligned between video and audio. Consequently, through the joint generation, we can expect higher alignment between the generated pair of data. While using a larger $\gamma$ may lead to more aligned loss distributions, it degrades the generation quality as we mentioned previously. In this paper, we set $\gamma$ to $1.5$, unless otherwise noted. How to automatically set this hyperparameter remains as future work.
III-C How to feed cross-modal features into U-Net
III-C1 The standard choice: cross-attention and its limitation
In the literature, the cross-attention mechanism has been extensively used for cross-modal conditioning in diffusion models. Figure 3 shows the simplest design in this approach adopted in our baseline method [5]. For brevity, we discuss the case of audio-conditioned video generation, but it can also be applied to the case of video-conditioned audio generation. In this design, the conditional audio information is embedded into a single feature vector by an encoder, and it is used to compute keys and values in the cross-attention taken with the intermediate features in the video generation model. By training the encoder and the cross-attention block with audio-video paired data, we can make the model generate videos aligned with given audio information. Although this design is simple and widely applicable, it is quite challenging to achieve higher temporal consistency between the audio condition and generated video, since the single vector does not have sufficient capability to represent every piece of temporally local information in the audio condition.
To boost the capability of the embedding features, we can extend the above-mentioned design by using multiple vectors each of which represents the temporally local information of conditional audio as done in [35]. However, we still cannot strongly guarantee the temporal alignment, as it provides too much flexibility to connect the temporally-local audio information with the video to be generated, which may cause mis-alignment. It is also possible to adopt a more sophisticated attention mechanism [4] or specifically dedicated encoder for embedding [25], but this substantially reduces applicability to the existing audio and video generation models, which does not fit our goal in this work.
@@ III-C1 The standard choice: cross-attention and its limitation
III-C2 Cross-Modal Conditioning as Positional Encoding (CMC-PE)
To achieve higher temporal alignment, we propose Cross-Modal Conditioning as Positional Encoding (CMC-PE), a simple but effective method of cross-modal conditioning. Figure 3 depicts how CMC-PE works. First, the conditional audio is encoded to a sequence of feature vectors along with time frames that work as if representing temporal position information. The extracted features are then added to the intermediate features in the video U-Net to function as positional embedding. To make this addition process valid, the features are interpolated and broadcast in advance to match their shape with that of the video features. Finally, the updated features are processed with a self-attention block. The features used for CMC-PE are extracted from current noisy latents $\mathbf{x}_{t}$ by the connector. We adopt the self-conditioning technique here [41], where the estimated data $\hat{\mathbf{x}}_{0|t}$ at each timestep shown in Eq. (7) is concatenated to the input.
@@ III-D Training and inference
where $s\in[\mathrm{v},\mathrm{a}]$ indicates the modality where the loss is to be computed, and $t_{\mathrm{v}}$ and $t_{\mathrm{a}}$ represent the local timestep for video and audio, respectively. A key point here is that the local timesteps are independently sampled from a uniform distribution. Due to this, the loss is computed on all possible combinations of the local timesteps so that the trained model can handle the timestep adjustment with any value of $\gamma$ specified in the inference phase. During the training of the model, the connectors and the inserted self-attention blocks are optimized with audio-video paired data, while the pre-trained parameters of the other modules are fixed, with one small exception (discussed in the appendix III-E).
@@
The generation process in our method is almost the same as that in usual diffusion models except for the setting of local timesteps. In each step of the generation process, the model predicts noise for video and audio latents following the local timestep setting. Once these noises are predicted, we can apply any inference technique, also called a “sampler” [1], to obtain $\mathbf{x}_{m(t-1)}^{(\mathrm{v})}$ and $\mathbf{x}_{n(t-1)}^{(\mathrm{a})}$ in the same manner as in the base models. In this paper, we use one of the most popular ones, DDIM [7] shown in Eq. (6), for both modalities. The generation process in the proposed method is summarized in Algorithm 1.
@@
III-E Implementation details